OpenStudy (anonymous):

The sum of the ages of Warren and Tatiana is 82 years. 11 years ago, Warren's age was 3 times Tatiana's age. How old is Warren now?

OpenStudy (anonymous):

Let warren' age=x then Tatiana's age=82-x 11 years back Warren'ns age==x-11 Tatiana's age=82-x-11=71-x according to the condition x-11=3(71-x) x-11=213-3x x+3x=213+11=224 4x=224 x=56 complete it.

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

this is easy